# FlightPatchNet: Multi-Scale Patch Network with Differential Coding for Flight Trajectory Prediction

### v1

@@ Abstract
Accurate multi-step flight trajectory prediction plays an important role in Air Traffic Control, which can ensure the safety of air transportation. Two main issues limit the flight trajectory prediction performance of existing works. The first issue is the negative impact on prediction accuracy caused by the significant differences in data range. The second issue is that real-world flight trajectories involve underlying temporal dependencies, and existing methods fail to reveal the hidden complex temporal variations and only extract features from one single time scale. To address the above issues, we propose FlightPatchNet, a multi-scale patch network with differential coding for flight trajectory prediction. Specifically, FlightPatchNet first utilizes the differential coding to encode the original values of longitude and latitude into first-order differences and generates embeddings for all variables at each time step. Then, a global temporal attention is introduced to explore the dependencies between different time steps. To fully explore the diverse temporal patterns in flight trajectories, a multi-scale patch network is delicately designed to serve as the backbone. The multi-scale patch network exploits stacked patch mixer blocks to capture inter- and intra-patch dependencies under different time scales, and further integrates multi-scale temporal features across different scales and variables. Finally, FlightPatchNet ensembles multiple predictors to make direct multi-step prediction. Extensive experiments on ADS-B datasets demonstrate that our model outperforms the competitive baselines. Code is available at: https://github.com/FlightTrajectoryResearch/FlightPatchNet.
@@ 1 Introduction
Flight Trajectory Prediction (FTP) is an essential task in the Air Traffic Control (ATC) procedure, which can be applied to various scenarios such as air traffic flow prediction Abadi6878453 ; LIN2019105113 , aircraft conflict detection AdeepGaussian , and arrival time estimation WANG2018280 . Accurate FTP can ensure the safety of air transportation and improve real-time airspace management LIN8846596 ; Shi9136843 . Generally, FTP tasks can be divided into three categories: long-term Jeong8190764 ; Runle7999617 , medium-term Yuan7554828 ; Chen2016ShortmediumtermPF , and short-term huang2017short ; duan2018unified . Among them, short-term trajectory prediction has the greatest impact on ATC and is increasingly in demand for air transportation. In this paper, we mainly focus on the short-term FTP task, which aims to predict future flight trajectories based on historical observations.
@@
In ATC domain, multi-step trajectory prediction can provide more practical applications than single-step prediction LIN8846596 . It can be divided into Iterated Multi-Step (IMS) prediction and Direct Multi-Step (DMS) prediction. IMS-based methods Yan6972562 ; Zhang2023FlightTP ; Guo2023FlightBERT make multi-step prediction recursively, which learns a single-step model and iteratively applies the predicted values as observations to forecast the next trajectory point. Due to the error accumulation problem and the step-by-step prediction scheme, this type of methods usually fails in multi-step prediction and cannot meet the real-time requirement. By contrast, DMS-based methods Guo2023FlightBERT++ ; wuhan2023bi can directly generate future trajectory points at once, which can tackle the error accumulation problem and improve prediction efficiency. Therefore, this paper performs the short-term FTP task in DMS way.
@@
However, two main issues are not well addressed in existing works Yan6972562 ; Zhang2023FlightTP ; Guo2023FlightBERT++ ; wuhan2023bi , limiting the trajectory prediction performance. The first issue is the negative impact on prediction accuracy caused by the significant differences in data range. In general, longitude and latitude are denoted by degree but altitude is by meter. Since one degree is approximately 111 kilometers, the data range of longitude and latitude are extremely different from that of altitude. Some previous works CNN-LSTM9145522 ; LSTM8489734 directly utilized normalization algorithms to scale variables into the same range, e.g., from 0 to 1. However, the actual prediction errors may be unacceptable for FTP task when evaluated in raw data range. FlightBERT Guo2023FlightBERT and FlightBERT++ Guo2023FlightBERT++ proposed binary encoding representation to convert variables from rounded decimal numbers to binary vectors, which regards the FTP task as multiple binary classification problem. Although BE representation can avoid the vulnerability caused by normalization algorithms, one serious limitation is introduced: a high bit misclassification in binary will lead to a large absolute error in decimal.
@@
The second issue is that real-world flight trajectories involve underlying temporal dependencies, and existing methods Shi9136843 ; Guo2023FlightBERT ; Guo2023FlightBERT++ fail to reveal the hidden complex temporal variations and only extract features from one single time scale. As shown in Figure 1, the original series of longitude and latitude are over-smoothing and obscure abundant temporal variations, which can be clearly observed from first-order difference series. Besides, the temporal variation patterns of longitude and latitude are quite distinct from those of altitude which have an obvious global trend but suffer from intense local fluctuations. For example, slight turbulence can exert a significant influence on the altitude but produce a negligible effect on the longitude and latitude. A single-scale model cannot simultaneously capture both local temporal details and global trends. This calls for powerful multi-scale temporal modeling capacity. Furthermore, if the learned multi-scale temporal patterns are simply aggregated, the model is failed to focus on contributed patterns. Meanwhile, it is essential to explore relationships across variables, e.g., the velocity at current time step directly effects the location at next time step. Thus, scale-wise correlations and inter-variable relationships should be fully considered when modeling the multi-scale temporal patterns.
Based on above analysis, this paper proposes a multi-scale patch network with differential coding (FlightPatchNet) to address above issues. Specifically, we utilize the differential coding to encode the original values of longitude and latitude into first-order differences and retain original values of other variables as inputs. Due to the dependencies between nearby and distant time steps, we introduce global temporal embedding to explore the correlations between time steps. Then, a multi-scale patch network is proposed to enable the ability of powerful and complete temporal modeling. The multi-scale patch network divide the trajectory series into patches of different sizes, and exploits stacked patch mixer blocks to capture global trends across patches and local details within patches. To further promote the multi-scale temporal modeling capacity, a multi-scale aggregator is introduced to capture scale-wise correlations and inter-variable relationships. Finally, FlightPatchNet ensembles multiple predictors to make direct multi-step forecasting, which can benefit from complementary multi-scale temporal features and improve the generalization ability. The main contributions are summarized as follows:
@@
  • We propose FlightPatchNet to fully explore underlying multi-scale temporal patterns. A multi-scale patch network is designed to capture inter- and intra-patch dependencies under different time scales, and integrate multi-scale temporal features across scales and variables. To our knowledge, this is the first work that introduces multi-scale modeling for flight trajectory prediction.
@@ 2 Related Work
Kinetics-and-Aerodynamics Methods
The Kinetics-and-Aerodynamics methods thipphavong2013adaptive ; soler2015multiphase ; benavides2014implementation ; tang20154d divide the entire flight process into several phases, and establish motion equations for each phase to formulate the flight status. For example, wang2009prediction adopted basic flight models to construct horizontal, vertical, and velocity profiles based on the characteristics of different flight phases. Zhijing7867472 combined the dynamics-and-kinematics models and grayscale theory to predict future trajectories. The grayscale theory can address the parameter missing problem in dynamics-and-kinematics models and improve the prediction performance. Due to numerous unknown and time-varying flight parameters of aircraft, these fixed-parameter methods cannot accurately describe the flight status, leading to a poor performance and limited application scenarios.
State-Estimation Methods
The Kalman Filter and its variants Yan6972562 ; wang20144d ; xi2008simulation are the typical single-model state-estimation algorithms for FTP task, which applies the predefined state equations to estimate the next flight status based on the current observation. For example, xi2008simulation applied the Kalman Filter to track discrete flight trajectories by calculating a continuous state transition matrix. However, single-model algorithms cannot adapt to the complex ATC environment. To address this issue, Interactive Multi Model algorithms hwang2003flight ; li2005survey have been proposed and successfully applied for trajectory analysis. Although multi-model algorithms can achieve better prediction performance, the computational complexity is high and cannot satisfy the real-time requirement.
Deep Learning Methods
With the rapid development of deep learning, there has been a surge of deep learning methods for FTP task Guo2023FlightBERT ; Guo2023FlightBERT++ ; pang2022bayesian ; xu2021multi ; Sahadevan . These learning-based approaches can extract high-dimensional features from raw data, which have achieved a more magnificent performance compared to previous methods. For example, Sahadevan used a Bi-directional Long-Short-Term-Memory (Bi-LSTM) network to explore both forward and backward dependencies in the sequential trajectory data. FlightBERT Guo2023FlightBERT employed binary encoding to represent the attributes of the trajectory points and considered the FTP task as a multi binary classification problem. However, these works predict the future trajectory recursively and suffer from serious error accumulation. Recently, FlightBERT++ Guo2023FlightBERT++ has been introduced for DMS prediction, which considers the prior horizon information and directly predicts the differential values between adjacent points.
@@ 3 Methodology
FTP task can be formulated as Multivariate Time Series (MTS) forecasting problem. Formally, given a sequence of historical observations $\mathbf{X}=\left\{\mathbf{x}_{1},...,\mathbf{x}_{L}\right\}\in\mathbb{R}^{C\times L}$, where $C$ is the state dimension, $L$ is the look-back window size and $\mathbf{x}_{t}\in\mathbb{R}^{C\times 1}$ denotes the flight state at time step $t$. The task is to predict future $T$ time steps $\hat{\mathbf{Y}}=\left\{\hat{\mathbf{x}}_{L+1},...,\hat{\mathbf{x}}_{L+T}\right\}\in\mathbb{R}^{C^{\prime}\times T}$, where $C^{\prime}$ is the predicted state dimension. Specifically, in this work, the flight state $\mathbf{x}_{t}$ represents longitude, latitude, altitude, and velocities along the previous three dimensions, i.e.,$\mathbf{x}_{t}=(Lon_{t},Lat_{t},Alt_{t},Vx_{t},Vy_{t},Vz_{t})^{\top}$.
Figure 2: FlightPatchNet architecture. (a) Global Temporal Embedding to explore the correlations between different time steps. (b) Multi-Scale Patch Network to capture inter- and intra-patch dependencies under different time scales and integrate temporal features across scales and variables. (c) Predictors to exploit complementary temporal features and make direct multi-step predictions.
The overall architecture of FlightPatchNet is shown in Figure 2, which consists of Global Temporal Embedding, Multi-Scale Patch Network, and Predictors. Global Temporal Embedding first utilizes differential coding to transform the original values of longitude and latitude into first-order differences and embeds all variables of the same time step into temporal tokens. A global temporal attention is then introduced to capture the inherent dependencies between different tokens. Multi-Scale Patch Network is proposed to serve as the backbone which is composed of stacked patch mixer blocks and a multi-scale aggregator. Stacked patch mixer blocks divide trajectory series into patches of different sizes from large scale to small scale. Based on divided patches, each patch mixer block exploits a patch encoder and decoder to capture inter- and intra-patch dependencies, endowing our model with powerful temporal modeling capability. To further integrate multi-scale temporal patterns, a multi-scale aggregator is incorporated into the network to capture scale-wise correlations and inter-variable relationships. Predictors provide direct multi-step trajectory forecasting and each predictor is a fully connected network. All the predictor results are aggregated to generate the final prediction.
@@ 3.1 Global Temporal Embedding
In the context of WGS84 Coordinate System, the longitude and latitude are limited to the intervals $[0,\pm 180^{\circ}]$ and $[0,\pm 90^{\circ}]$ respectively, while the altitude can span from 0 up to tens of thousands of meters. The significant differences of data range caused by physical units may impair the trajectory prediction performance. Generally, normalization algorithms are applied to address this issue. However, the normalized prediction errors should be transformed into raw data range to evaluate the actual performance. For example, if the absolute prediction error of longitude is $10^{-4}$ after using Min-Max normalization algorithm, the actual prediction error is 0.036° (approximate 4000 meters), which is unacceptable for FTP task. Moreover, as shown in Figure 1, the original series of longitude and latitude are over-smoothing and only reflect the overall flight trend over a period. If temporal patterns are learned from the original values of longitude and latitude, the model is failed to explore the implicit semantic information and cannot focus on short-term temporal variations in flight trajectories.
@@
[…]
@@
Given the trajectory series $\mathbf{X}\in\mathbb{R}^{C\times L}$, we first project flight state at each time step into $d$ dimension to generate temporal embeddings $\mathbf{T}^{0}\in\mathbb{R}^{L\times d}$. Then, we apply multi-head self-attention (MSA) Vaswani2017AttentionIA on the dimension $L$ to capture the dependencies across all time steps. After attention, the embedding at each time step is enriched with temporal information from other time steps. This process is formulated as:
[…]
where $l$ is the number of attention layers, $LayerNorm$ denotes the layer normalization ba2016layer which has been widely adopted to address non-stationary issues, $MSA$ is the multi-head self-attention layer, $FC$ denotes a fully-connected layer and $Linear$ projects the embedding of each time step to dimension $C$, i.e., $\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$.

### v3

@@ Abstract
Accurate multi-step flight trajectory prediction plays an important role in Air Traffic Control, which can ensure the safety of air transportation. Two main issues limit the flight trajectory prediction performance of existing works. The first issue is the negative impact on prediction accuracy caused by the significant differences in data range. The second issue is that real-world flight trajectories involve underlying temporal dependencies, and most existing methods fail to reveal the hidden complex temporal variations and extract features from one single time scale. To address the above issues, we propose FlightPatchNet, a multi-scale patch network with differential coding for flight trajectory prediction. Specifically, FlightPatchNet first utilizes differential coding to encode the original values of longitude and latitude into first-order differences and generates embeddings for all variables at each time step. Then, global temporal attention is introduced to explore the dependencies between different time steps. To fully explore the diverse temporal patterns in flight trajectories, a multi-scale patch network is delicately designed to serve as the backbone. The multi-scale patch network exploits stacked patch mixer blocks to capture inter- and intra-patch dependencies under different time scales, and further integrates multi-scale temporal features across different scales and variables. Finally, FlightPatchNet ensembles multiple predictors to make direct multi-step prediction. Extensive experiments on ADS-B datasets demonstrate that our model outperforms the competitive baselines.
@@ 1 Introduction
Flight Trajectory Prediction (FTP) is an essential task in the Air Traffic Control (ATC) procedure, which can be applied to various scenarios such as air traffic flow prediction [Abadi et al., 2015, Lin et al., 2019], aircraft conflict detection [Chen et al., 2020], and arrival time estimation [Wang et al., 2018]. Accurate FTP can ensure the safety of air transportation and improve real-time airspace management [Lin et al., 2020, Shi et al., 2021]. Generally, FTP tasks can be divided into three categories: long-term [Jeong et al., 2017, Runle et al., 2017], medium-term [Yuan et al., 2016, Chen et al., 2016], and short-term [Huang et al., 2017, Duan et al., 2018]. Among them, short-term trajectory prediction has the greatest impact on ATC and is increasingly in demand for air transportation. In this paper, we mainly focus on the short-term FTP task, which aims to predict future flight trajectories based on historical observations.
@@
In the ATC domain, multi-step trajectory prediction can provide more practical applications than single-step prediction [Lin et al., 2020]. It can be divided into Iterated Multi-Step (IMS) prediction and Direct Multi-Step (DMS) prediction. IMS-based methods [Yan et al., 2013, Zhang et al., 2023, Guo et al., 2023] make multi-step prediction recursively, which learns a single-step model and iteratively applies the predicted values as observations to forecast the next trajectory point. Due to the error accumulation problem and the step-by-step prediction scheme, this type of methods usually fails in multi-step prediction and has poor real-time performance. By contrast, DMS-based methods [Wu et al., 2023b, Guo et al., 2024] can directly generate future trajectory points at once, which can tackle the error accumulation problem and improve prediction efficiency. Therefore, this paper performs the short-term FTP task in DMS way.
@@
However, two main issues are not well addressed in existing works [Yan et al., 2013, Zhang et al., 2023, Wu et al., 2023b, Guo et al., 2024], limiting the trajectory prediction performance. The first issue is the negative impact on prediction accuracy caused by the significant differences in data range. In general, longitude and latitude are denoted by degree but altitude is by meter. Since one degree is approximately 111 kilometers, the data range of longitude and latitude are extremely different from that of altitude. Some previous works [Shi et al., 2018, Ma and Tian, 2020] directly utilized normalization algorithms to scale variables into the same range, e.g., from 0 to 1. However, the actual prediction errors could be large for FTP tasks when evaluated in raw data range (as shown in Table 2). FlightBERT [Guo et al., 2023] and FlightBERT++ [Guo et al., 2024] proposed binary encoding (BE) representation to convert variables from rounded decimal numbers to binary vectors, which regards the FTP task as multiple binary classification problem. Although BE representation can avoid the vulnerability caused by normalization algorithms, one serious limitation is introduced: a high bit misclassification in binary will lead to a large absolute error in decimal.
@@
The second issue is that real-world flight trajectories involve underlying temporal dependencies, and most existing methods [Shi et al., 2021, Guo et al., 2023, 2024] fail to reveal the hidden complex temporal variations and extract features from one single time scale. As shown in Figure 1, the original series of longitude and latitude are over-smoothing and obscure abundant temporal variations, which can be observed from the first-order difference series. Besides, the temporal variation patterns of longitude and latitude are quite distinct from those of altitude which have an obvious global trend but suffer from intense local fluctuations. For example, slight turbulence can exert a significant influence on the altitude but produce a negligible effect on the longitude and latitude. A single-scale model cannot simultaneously capture both local temporal details and global trends [Wu et al., 2023a, Wang et al., 2023]. This calls for powerful multi-scale temporal modeling capacity. Furthermore, if the learned multi-scale temporal patterns are simply aggregated, the model is failed to focus on contributed patterns [Chen et al., 2023a]. Meanwhile, it is essential to explore relationships across variables [Zhang and Yan, 2023, Han et al., 2024], e.g., the velocity at current time step directly affects the location at next time step. Thus, scale-wise correlations and inter-variable relationships should be fully considered when modeling the multi-scale temporal patterns.
Based on above analysis, this paper proposes a multi-scale patch network with differential coding (FlightPatchNet) to address above issues. Specifically, we utilize differential coding to encode the original values of longitude and latitude into first-order differences and retain the original values of other variables as inputs. Due to the dependencies between nearby and distant time steps, we introduce global temporal embedding to explore the correlations between time steps. Then, a multi-scale patch network is proposed to enable the ability of powerful and complete temporal modeling. The multi-scale patch network divides the trajectory series into patches of different sizes, and exploits stacked patch mixer blocks to capture global trends across patches and local details within patches. To further promote the multi-scale temporal modeling capacity, a multi-scale aggregator is introduced to capture scale-wise correlations and inter-variable relationships. Finally, FlightPatchNet ensembles multiple predictors to make direct multi-step forecasting, which can benefit from complementary multi-scale temporal features and improve the generalization ability. The main contributions are summarized as follows:
@@
  • We propose FlightPatchNet to fully explore underlying multi-scale temporal patterns. A multi-scale patch network is designed to capture inter- and intra-patch dependencies under different time scales, and integrate multi-scale temporal features across scales and variables.
@@ 2 Related Work
Kinetics-and-Aerodynamics Methods The Kinetics-and-Aerodynamics methods [Thipphavong et al., 2013, Benavides et al., 2014, Soler et al., 2015, Tang et al., 2015] divide the entire flight process into several phases, and establish motion equations for each phase to formulate the flight status. For example, Wang et al. [2009] adopted basic flight models to construct horizontal, vertical, and velocity profiles based on the characteristics of different flight phases. Zhou et al. [2016] combined the dynamics-and-kinematics models and grayscale theory to predict future trajectories. The grayscale theory can address the parameter missing problem in dynamics-and-kinematics models and improve the prediction performance. Due to numerous unknown and time-varying flight parameters of aircraft, these fixed-parameter methods cannot accurately describe the flight status, leading to poor performance and limited application scenarios.
State-Estimation Methods The Kalman Filter and its variants [Xi et al., 2008, Yan et al., 2013] are the typical single-model state-estimation algorithms for FTP tasks, which applies the predefined state equations to estimate the next flight status based on the current observation. For example, Xi et al. [2008] applied the Kalman Filter to track discrete flight trajectories by calculating a continuous state transition matrix. However, single-model algorithms cannot adapt to the complex ATC environment. To address this issue, Interactive Multi-Model algorithms [Hwang et al., 2003, Li and Jilkov, 2005] have been proposed and successfully applied for trajectory analysis. Although multi-model algorithms can achieve better prediction performance, the computational complexity is high and cannot satisfy the real-time requirement.
Deep Learning Methods With the rapid development of deep learning, there has been a surge of deep learning methods for FTP task [Xu et al., 2021, Pang et al., 2022, Sahadevan et al., 2022, Zhang et al., 2023, Guo et al., 2023, 2024]. These learning-based approaches can extract high-dimensional features from raw data, which have achieved a more magnificent performance compared to previous methods. For example, Sahadevan et al. [2022] used a Bi-directional Long-Short-Term-Memory (Bi-LSTM) network to explore both forward and backward dependencies in the sequential trajectory data. Zhang et al. [2023] proposed a wavelet transform-based framework (WTFTP) to perform time-frequency analysis of flight patterns for trajectory prediction. FlightBERT [Guo et al., 2023] employed binary encoding to represent the attributes of the trajectory points and considered the FTP task as a multi-binary classification problem. However, these works predict the future trajectory recursively and suffer from serious error accumulation. Recently, FlightBERT++ [Guo et al., 2024] has been introduced for DMS prediction, which considers the prior horizon information and directly predicts the differential values between adjacent points.
@@ 3 Methodology
The short-term FTP task can be formulated as a Multivariate Time Series (MTS) forecasting problem. Given a sequence of historical observations $\mathbf{X}=\left\{\mathbf{x}_{1},...,\mathbf{x}_{L}\right\}\in\mathbb{R}^{C\times L}$, where $C$ is the state dimension, $L$ is the look-back window size and $\mathbf{x}_{t}\in\mathbb{R}^{C\times 1}$ is the flight state at time step $t$. The task is to predict future $T$ time steps $\hat{\mathbf{Y}}=\left\{\hat{\mathbf{x}}_{L+1},...,\hat{\mathbf{x}}_{L+T}\right\}\in\mathbb{R}^{C^{\prime}\times T}$, where $C^{\prime}$ is the predicted state dimension. In this work, the flight state $\mathbf{x}_{t}$ represents longitude, latitude, altitude, and velocities along the previous three dimensions, i.e., $\mathbf{x}_{t}=(Lon_{t},Lat_{t},Alt_{t},Vx_{t},Vy_{t},Vz_{t})^{\top}$.
The overall architecture of FlightPatchNet is shown in Figure 2, which consists of Global Temporal Embedding, Multi-Scale Patch Network, and Predictors. Global Temporal Embedding first utilizes differential coding to transform the original values of longitude and latitude into first-order differences and embeds all variables of the same time step into temporal tokens. Global temporal attention is then introduced to capture the inherent dependencies between different tokens. Multi-Scale Patch Network is proposed to serve as the backbone which is composed of stacked patch mixer blocks and a multi-scale aggregator. Stacked patch mixer blocks divide trajectory series into patches of different sizes from large scale to small scale. Based on divided patches, each patch mixer block exploits a patch encoder and decoder to capture inter- and intra-patch dependencies, endowing our model with powerful temporal modeling capability. To further integrate multi-scale temporal patterns, a multi-scale aggregator is incorporated into the network to capture scale-wise correlations and inter-variable relationships. Predictors provide direct multi-step trajectory forecasting and each predictor is a fully connected network. All the predictor results are aggregated to reconstruct the final prediction trajectory.
Figure 2: FlightPatchNet architecture. (a) Global Temporal Embedding to explore the correlations between different time steps. (b) Multi-Scale Patch Network to capture inter- and intra-patch dependencies under different time scales and integrate temporal features across scales and variables. (c) Predictors to exploit complementary temporal features and make direct multi-step prediction.
@@ 3.1 Global Temporal Embedding
In the context of the WGS84 Coordinate System, the longitude and latitude are limited to the intervals $180^{\circ},180^{\circ}]$ and $90^{\circ},90^{\circ}]$ respectively, while the altitude can span from 0 up to tens of thousands of meters. The significant differences of data range caused by physical units may impair the trajectory prediction performance. Generally, normalization algorithms are applied to address this issue. However, the normalized prediction errors should be transformed into raw data range to evaluate the actual performance. For example, if the absolute prediction error of longitude is $10^{-4}$ after using the Min-Max normalization algorithm, the actual prediction error is 0.036° (approximately 4000 meters). Moreover, as shown in Figure 1, the original series of longitude and latitude are over-smoothing and reflect the overall flight trend over a period. If temporal patterns are learned from the original values of longitude and latitude, the model fails to explore the implicit semantic information and cannot focus on short-term temporal variations in flight trajectories.
@@
where $\phi$ denotes the longitude, $\varphi$ denotes the latitude, and $R$ is the radius of the earth. By using differential coding for longitude and latitude, the differences in data range are effectively reduced. For example, in our dataset, the range of latitude in original data is about $46^{\circ},70^{\circ}]$ and that in differential data is about [3860m,3860m][-3860m,3860m], which spans a similar data range as the altitude. Compared to the original sequences, the differential series can explicitly reflect the underlying temporal variations, which is essential for short-term temporal modeling. Besides, adopting the first-order differences instead of second- or higher-order differences enables the model to reconstruct the predicted trajectory based on the last observation. Note that we utilize the original values of altitude as inputs rather than differential values. One important reason is that altitude is more susceptible to noise, failing to reflect the actual temporal variations. To this end, the flight state at time step $t$ becomes $\mathbf{x}_{t}=(\Delta_{\mathit{Lon}},\Delta_{\mathit{Lat}},Alt_{t},Vx_{t},Vy_{t},Vz_{t})^{\top}$.
@@
Given the trajectory series $\mathbf{X}\in\mathbb{R}^{C\times L}$, we first project flight state at each time step into $d$ dimension to generate temporal embeddings $\mathbf{T}^{0}\in\mathbb{R}^{L\times d}$. Then, we apply multi-head self-attention (MSA) [Vaswani et al., 2017] on the dimension $L$ to capture the dependencies across all time steps. After attention, the embedding at each time step is enriched with temporal information from other time steps. This process is formulated as:
$\begin{aligned} \mathbf{T}^{0}&=\mathit{TimeEmbedding}(\mathbf{X}^{\top})\\ \mathbf{T}^{i}&=\mathit{LayerNorm}(\mathbf{T}^{i-1}+\mathit{MSA}(\mathbf{T}^{i-1}),i=1,\dots,l\\ \mathbf{T}^{i}&=\mathit{LayerNorm}(\mathbf{T}^{i}+FC(\mathbf{T}^{i}),i=1,\dots,l\\ \mathbf{Z}&={(\mathit{Linear}(\mathbf{T}^{l}}))^{\top}\end{aligned}$
where $l$ is the number of attention layers, $LayerNorm$ denotes the layer normalization [Ba et al., 2016] which has been widely adopted to address non-stationary issues, $MSA$ is the multi-head self-attention layer, $FC$ denotes a fully-connected layer and $Linear$ projects the embedding of each time step to dimension $C$, i.e., $\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$.